An alloy of Cu,Ag &Au is found to have copper constituting the FCC lattice.If silver atoms occupy the edge centres and gold is present at body centre,the alloy has the formula?
 

Grade:12th pass

1 Answers

Arun
5 years ago
Effective no. of Cu atoms in a unit cell = 8 * 1/8 + 6 * ½ = 1 + 3 = 4
 
Effective No. of Ag atoms in a unit cell = 12 * ¼ = 3
Effective no. of Au atoms in a unit cell = 1
So, general formula of compound is Cu4Ag3Au